# Orbital optical activity in noncentrosymmetric metals and superconductors

### v1

@@ Abstract
We present the optical activity induced by the orbital magnetic moment in metals and superconductors using Green’s function formalization. We show that an apparent singularity of the optical activity vanishes in the normal state; however, it remains finite in the superconducting state and is related to the superconducting Edelstein effect, ensuring the missing area measurement. Finally, we calculate the optical activity in a model Hamiltonian mimicking doped transition metal dichalcogenides to investigate its characteristic spectrum, and we analyze the Kerr effect to discuss a possibility to observe the optical activity in experiments.
@@ I Introduction
In this paper, we discuss the optical activity originating from orbital magnetic moments in metals and superconductors using a Green’s function formalism. We discuss a no-go theorem concerning a singularity of the optical activity in Secs. II and III, and we show that the singularity appearing in the superconducting state is related to the superconducting Edelstein effect. Furthermore, we present model calculations investigating two-dimensional materials such as transition metal dichalcogenides and twisted bilayer graphene in the normal state and the superconducting state in Secs. IV and V, respectively. We also analyze the Kerr spectrum. Finally, in Sec. VI, we conclude this paper and discuss the possibility of an experimental observation.
@@
In this section, we derive the orbital optical activity in the normal state using a Green’s function method. The optical activity originates from the wavenumber dependence of the optical conductivity $\sigma(\bm{q},\Omega)$, where $\bm{q}$ and $\Omega$ are the wavenumber and the frequency of the irradiation light. $\sigma(\bm{q}=0,\Omega)$ represents the uniform conductivity that changes the refractive index of a material, leading to observable effects such as the reflectance and the rotation of linearly-polarized light (Faraday rotation and Kerr rotation) in magnets. In noncentrosymmetric systems, the first-order term in the wavenumber of the optical conductivity $\bm{\partial}_{q}\sigma(\bm{q}=0,\Omega)\cdot\bm{q}$ can be significant. This is the optical activity, which is the main quantity in this paper, exhibiting optical rotation, dichroism, and birefringence. It includes two parts: a spin part and an orbital part. In this paper, we only focus on the orbital part.
@@ II.1 Green’s function formalism of the orbital optical activity
Here, we use $\partial_{\lambda}=\partial/\partial k_{\lambda}$. We note the identity $\partial_{\lambda}G^{R/A}=G^{R/A}v_{\bm{k}}^{\lambda}G^{R/A}$, which is useful for numerical calculations. Finally, we obtain the optical activity $\sigma_{\mu\nu\lambda}(\Omega)=\Phi_{\mu\nu\lambda}(\Omega)/i(\Omega+i\delta)$, where $\delta=+0$ is an adiabatic factor. This tensor $\sigma_{\mu\nu\lambda}$ describes all optical activity phenomena induced by the orbital moment, including the NOA and the OMEE. This formulation has been useful for analyses in periodic crystals as it avoids difficulties such as origin dependence and gauge invariance. Some works have derived this tensor with the band representation and proved explicitly that this formalism avoids these difficulties, providing a physical interpretation of the obtained gauge-invariant equations [21, 23, 41, 42, 27].
@@ II.2 No-go theorem
The optical activity has a singularity at $\Omega=0$ because of the factor $1/(\Omega+i\delta)=1/\Omega-i\pi\delta(\Omega)$. This $\delta$-function and the accompanying anomalous divergence scaling $\Omega^{-1}$ are known as the false divergences problem when using the velocity gauge from the minimal-coupling Hamiltonian [34]. This divergence is supposed to vanish in the normal state, which has been shown using the Bloch band representation [27]. We here clarify whether it also vanishes using the Green’s function formula. The coefficient in front of the $\delta$ function in the optical activity reads
@@
Here, $\varepsilon_{\mu\nu\lambda}$ is the totally antisymmetric tensor. Almost all components of $\Phi_{\mu\nu\lambda}(0)$ are zero. Only totally antisymmetric terms appear due to $\alpha$, which is real (see Appendix A for a detailed derivation). This term corresponds to the chiral magnetic effect, where a current density is induced by a magnetic field as $\bm{J}=\alpha\bm{B}$. However, this phenomenon usually does not occur unless, for example, the chiral chemical potential $\mu_{5}$ is finite in the non-equilibrium case [43]. Using the Bloch band representation, this $\alpha$ is shown to be zero, resulting from its topological nature [23, 27]. However, the formula using Green’s function can not immediately prove that this term vanishes. Therefore, in the following discussion, we only consider cases where the chiral magnetic effect vanishes due to crystal point group symmetry.
@@ III.1 Green’s function formula for a superconductor
$H^{\mathrm{BdG}}_{\bm{k}nm}$ $=\begin{pmatrix}H_{\bm{k}}-\mu&-\Delta_{\bm{k}}\\ -\Delta^{\dagger}_{\bm{k}}&-H^{\mathrm{T}}_{-\bm{k}}+\mu\end{pmatrix}.$ (7d)
@@ III.2 Singularity and relation to the superconducting Edelstein effect
In the previous section, we have shown that the optical activity appears to have a $\delta$-function singularity at $\Omega=0$, which, however, vanishes in the normal state. In the superconducting case, the singularity generally is finite. The absence of cancellation is attributed to the difference in the velocity operator from the normal state. The coefficient of the $\delta$ function satisfies (see Appendix A for a detailed derivation)
@@
Here, $\mathcal{K}_{\mu\nu}$ is real because $\Phi_{\mu\nu\lambda}(0)$ is purely imaginary. Substituting Eq. (13a) into the original linear response relation, we obtain
@@
$\mathcal{K}$ represents the superconducting Edelstein effect (SEE), where a supercurrent induces a magnetization $\bm{M}$ or the magnetic field $\bm{B}$ induces the supercurrent. Thus, $\Phi_{\mu\nu\lambda}(0)$ also has the physical meaning of the SEE for the orbital contribution, in analogy to the case of the spin contribution.
@@
Here, we propose another means of observing this SEE. By measuring the spectrum of the optical activity, we can obtain $\Phi_{\mu\nu\lambda}(0)$, which is equivalent to the SEE. $\Phi_{\mu\nu\lambda}(0)$ can be obtained by measuring the coefficient of the anomalous divergence of $\mathrm{Re}\sigma_{\mu\nu\lambda}(\Omega)$ scaling as $\Omega^{-1}$. Furthermore, the spectrum of $\mathrm{Im}\sigma_{\mu\nu\lambda}(\Omega)$ can also determine the value of $\Phi_{\mu\nu\lambda}(0)$. Reference [40] has shown that the missing area, the area reduced from the spectrum of the normal phase created by the superconducting gap, is identical to $\Phi_{\mu\nu\lambda}(0)$ because of a universal sum rule and the no-go theorem in the case of the spin contribution. This method is useful for dirty superconductors, which is a common situation. As discussed in Sec. II.2, the no-go theorem is also established for the orbital contribution. Thus, the missing area measurement is also valid for the orbital contribution.
@@ IV.1 Optical activity in the normal state
$H_{\bm{k}}=\begin{pmatrix}G&\epsilon_{\bm{k}}\\ \epsilon^{*}_{\bm{k}}&-G\end{pmatrix}\otimes\sigma_{0},$ (15c)
$\epsilon_{\bm{k}}=te^{i\bm{k}\cdot\bm{a}_{1}}+t^{\prime}(e^{i\bm{k}\cdot\bm{a}_{2}}+e^{i\bm{k}\cdot\bm{a}_{3}}).$ (15d)
@@
$\int_{\mathrm{BZ}}\frac{d^{d}k}{(2\pi)^{d}}\sum_{n}f(\epsilon_{n\bm{k}}-\mu)\partial_{k_{i}}m^{\mathrm{orb}}_{n\bm{k}},$ (16)
where $m^{\mathrm{orb}}_{n\bm{k}}$ is the orbital magnetic moment oriented in the direction perpendicular to the surface. Without stretching $\mathrm{MoS_{2}}$, orbital magnetic moments (OMMs) with opposite signs exist around the $\mathrm{K}$ and $\mathrm{K}^{\prime}$ points [59]; however, the OMMD vanishes due to $\mathrm{C_{3v}}$ symmetry. When stretching $\mathrm{MoS_{2}}$, the OMMD in the same direction as the stretching direction becomes finite and generates the GME. This discussion is similar to the Berry curvature dipole in the nonlinear Hall effect [60]. The GME in the DC limit has been experimentally observed as the orbital Edelstein effect in stretched $\mathrm{MoS_{2}}$ [61, 62]. Above the frequency corresponding to the optical gap $\sim 2\mu$, interband effects become finite around $\mathrm{K}$ and $\mathrm{K}^{\prime}$ points.
@@
This peak shift is unique to the orbital optical activity in metals. It becomes clearer when the optical activity is represented by the Bloch band basis. The full representation is derived in a recent work [27] (see Eq. (26) in the paper). It shows that there is a Fermi surface effect affecting the interband transition besides the Fermi sea effect. This is unique to the orbital contribution. It demonstrates that the effect of optical transitions in the optical activity cannot be simply described by transitions from occupied to unoccupied states, which is described by the Fermi sea term. We note that in this two-band model with particle-hole symmetry, the Fermi sea term is suppressed. The Fermi surface term is locally enhanced at the optical gap $2\mu$, which is consistent with the result in Fig. 2.
IV.2 Optical Kerr effect
The optical activity $\sigma_{xyx}$ in our model can be observed using the optical Kerr effect. The Kerr effect is usually caused by the optical Hall effect in magnets without time-reversal symmetry. Because the present model obeys time-reversal symmetry, the Kerr effect originating from the optical Hall effect does not occur. However, the optical activity also generates an optical Kerr effect by a different mechanism, as described in the following discussion.
The spectrum of the longitudinal optical conductivity $\sigma_{ii}(\omega)$ can be extracted from the experimentally observed complex reflectance. Similarly, the spectrum of the optical activity can be extracted from the experimentally observed rotational angle and the ellipticity caused by the Kerr effect or the Faraday effect. In two-dimensional materials, transmission through the material does not exist, and thus, only the Kerr effect is allowed. In this subsection, we discuss the Kerr effect as a possibility of observing the optical activity experimentally.
@@ V Model calculation in the superconducting state
Figure 5: Numerical results of the Kerr rotation $\theta$ and ellipticity $\varepsilon$ for the reflected light (Left) and the transmitted light (Right) on a logarithmic scale. We denote the superconducting state and the normal state as sc and n, respectively. We use the same parameters and units as in Fig. 1. We set $\Delta=0.005$.
The Kerr rotation and ellipticity are also changed because of the spectral modification by the superconducting state. We show the Kerr rotation $\theta$ and ellipticity $\varepsilon$ in Fig. 5 (blue lines). We also show the results of the normal state for comparison (black lines). As discussed in Sec. IV.2, in the normal state, Eqs. (18) and (19) are almost purely imaginary in the Drude regime. Thus, the phase difference of the electric fields is almost $\pi/2$, and the ellipticity is dominant. In the superconducting state, for frequencies inside the superconducting gap, the dominant components of $\sigma_{xx/yy}$ and $\sigma_{xyx}$ are changed to the imaginary part and the real part, respectively. Therefore, Eq. (18) is no longer purely imaginary, and the Kerr rotation is enhanced. At low frequencies, $\mathrm{Im}\sigma_{xx}$ increases by $\Omega^{-1}$ and becomes comparable to the inverse vacuum impedance $(\mu_{0}c)^{-1}$. Then, the real part and imaginary part of Eq. (18) are comparable, and the rotational angle and ellipticity are in the same order of magnitude for the reflected light. For transmitted light, Eq. (19) is also no longer purely imaginary and becomes rather real when $(\mu_{0}c)^{-1}>\mathrm{Im}\sigma_{xx}$, thus the rotational angle is dominant. As the frequency decreases further, $\mathrm{Im}\sigma_{xx}$ increases by $\Omega^{-1}$ and becomes comparable to $(\mu_{0}c)^{-1}$. Then, the rotational angle and the ellipticity also become comparable in magnitude.

### v2

@@ Abstract
We present the optical activity induced by the orbital magnetic moment in metals and superconductors using Green’s function formalization. Usually, an apparent singularity of the optical activity vanishes in the normal state; however, we show that it remains finite in the superconducting state and is related to the superconducting Edelstein effect, ensuring the missing area measurement. Finally, we calculate the optical activity in a model Hamiltonian mimicking doped transition metal dichalcogenides to investigate its characteristic spectrum, which reveals a peak shift caused by the interband Fermi surface, its relation with the orbital magnetic moment, and the change of the spectrum in a superconductor at low frequency. We also analyze the optical rotation to discuss the possibility of observing the optical activity in experiments.
@@ I Introduction
In this paper, we discuss the optical activity originating from orbital magnetic moments in metals and superconductors using a Green’s function formalism. We discuss a no-go theorem concerning a singularity of the optical activity in Secs. II and III, and we show that the singularity appearing in the superconducting state is related to the superconducting Edelstein effect. Furthermore, we present model calculations investigating two-dimensional materials such as transition metal dichalcogenides and twisted bilayer graphene in the normal state and the superconducting state in Secs. IV and V, respectively. We also analyze the spectrum of the optical rotation. Finally, in Sec. VI.1, we conclude this paper and discuss the possibility of an experimental observation.
@@
In this section, we derive the orbital optical activity in the normal state using a Green’s function method. The optical activity originates from the wavenumber dependence of the optical conductivity $\sigma(\bm{q},\Omega)$, where $\bm{q}$ and $\Omega$ are the wavenumber and the frequency of the irradiation light. $\sigma(\bm{q}=0,\Omega)$ represents the uniform conductivity that changes the refractive index of a material, leading to observable effects such as the reflectance and the rotation of linearly-polarized light in magnets. In noncentrosymmetric systems, the first-order term in the wavenumber of the optical conductivity $\bm{\partial}_{q}\sigma(\bm{q}=0,\Omega)\cdot\bm{q}$ can be significant. This is the optical activity, which is the main quantity in this paper, exhibiting optical rotation, dichroism, and birefringence. It includes two parts: a spin part and an orbital part. In this paper, we only focus on the orbital part.
@@ II.1 Green’s function formalism of the orbital optical activity
Here, we use $\partial_{\lambda}=\partial/\partial k_{\lambda}$. We note the identity $\partial_{\lambda}G^{R/A}=G^{R/A}v_{\bm{k}}^{\lambda}G^{R/A}$, which is useful for numerical calculations. Finally, we obtain the optical activity $\sigma_{\mu\nu\lambda}(\Omega)=\Phi_{\mu\nu\lambda}(\Omega)/i(\Omega+i\delta)$, where $\delta=+0$ is an adiabatic factor. This tensor $\sigma_{\mu\nu\lambda}$ describes all optical activity phenomena induced by the orbital moment, including the NOA (antisymmetric part by interchange of indices $\mu$ and $\nu$) and the OMEE (symmetric part). This formulation has been useful for analyses in periodic crystals as it avoids difficulties such as origin dependence and gauge invariance. Some works have derived this tensor with the band representation and proved explicitly that this formalism avoids these difficulties, providing a physical interpretation of the obtained gauge-invariant equations [21, 23, 41, 42, 27].
@@ II.2 No-go theorem
The optical activity has a singularity at $\Omega=0$ because of the factor $1/(\Omega+i\delta)=1/\Omega-i\pi\delta(\Omega)$. This $\delta$-function and the accompanying anomalous divergence scaling $\Omega^{-1}$ are known as the false divergences problem when using the velocity gauge from the minimal-coupling Hamiltonian [34]. This divergence is supposed to vanish in the normal state (no-go theorem), which has been shown using the Bloch band representation [27]. We can also reconfirm that the singularity vanishes for almost all of the components of $\sigma_{\mu\nu\lambda}$ using the Green’s function formula as seen in the following.
The coefficient in front of the $\delta$ function in the optical activity reads
@@
Here, $\varepsilon_{\mu\nu\lambda}$ is the totally antisymmetric tensor. Almost all components of $\Phi_{\mu\nu\lambda}(0)$ are zero. Only totally antisymmetric terms appear due to $\alpha$, which is real (see Appendix A for a detailed derivation). This term corresponds to the chiral magnetic effect, where a current density is induced by a magnetic field as $\bm{J}=\alpha\bm{B}$. However, this phenomenon usually does not occur unless, for example, the chiral chemical potential $\mu_{5}$ is finite in the non-equilibrium case [43]. Using the Bloch band representation, this $\alpha$ is shown to be zero, resulting from its topological nature [23, 27]. However, the formula using Green’s function can not immediately prove that this term vanishes. In conclusion, the no-go theorem can be checked for all components except for the totally-antisymmetric term. We leave the general proof whether this term vanishes using Green’s functions assuming, e.g., strongly correlated systems as a future work. Thus, in the following discussion, we only consider cases where the totally-antisymmetric component vanishes due to crystal point group symmetry.
@@ III.1 Green’s function formula for a superconductor
$H^{\mathrm{BdG}}_{\bm{k}}$ $=\begin{pmatrix}H_{\bm{k}}-\mu&-\Delta_{\bm{k}}\\ -\Delta^{\dagger}_{\bm{k}}&-H^{\mathrm{T}}_{-\bm{k}}+\mu\end{pmatrix}.$ (7b)
@@ III.2 Singularity and relation to the superconducting Edelstein effect
In the previous section, we have seen that the optical activity appears to have a $\delta$-function singularity at $\Omega=0$, which, however, vanishes in the normal state. In the superconducting case, the singularity generally is finite. The absence of cancellation is attributed to the difference in the velocity operator from the normal state. The coefficient of the $\delta$ function satisfies (see Appendix A for a detailed derivation)
@@
Here, $\mathcal{K}_{\mu\nu}$ is real because $\Phi_{\mu\nu\lambda}(0)$ is purely imaginary. Substituting Eq. (13a) into the original linear response relation ($J_{\mu}=\Phi_{\mu\nu}(\bm{q},\Omega)A_{\nu}$), we obtain
@@
This equation shows that $\mathcal{K}$ represents the superconducting Edelstein effect (SEE), where a supercurrent induces a magnetization $\bm{M}$ or the magnetic field $\bm{B}$ induces the supercurrent. Thus, $\Phi_{\mu\nu\lambda}(0)$ also has the physical meaning of the SEE for the orbital contribution, in analogy to the case of the spin contribution.
@@
Here, we propose another means of observing this SEE. By measuring the spectrum of the optical activity, we can obtain $\Phi_{\mu\nu\lambda}(0)$, which is equivalent to the SEE. $\Phi_{\mu\nu\lambda}(0)$ can be obtained by measuring the coefficient of the anomalous divergence of $\mathrm{Re}\sigma_{\mu\nu\lambda}(\Omega)$ scaling as $\Omega^{-1}$. Furthermore, the spectrum of $\mathrm{Im}\sigma_{\mu\nu\lambda}(\Omega)$ can also determine the value of $\Phi_{\mu\nu\lambda}(0)$. Reference [40] has shown that the missing area, the area reduced from the spectrum of the normal phase created by the superconducting gap, is identical to $\Phi_{\mu\nu\lambda}(0)$ because of a universal sum rule and the no-go theorem in the case of the spin contribution. This method is useful for dirty superconductors, which is a common situation. As discussed in Sec. II.2 and this section, the no-go theorem is also established for the orbital contribution. Thus, the missing area measurement is also valid for the orbital contribution.
@@ IV.1 Optical activity in the normal state
$H_{\bm{k}}=\begin{pmatrix}G&\epsilon_{\bm{k}}\\ \epsilon^{*}_{\bm{k}}&-G\end{pmatrix}\otimes\sigma_{0},$ (15a)
$\epsilon_{\bm{k}}=te^{i\bm{k}\cdot\bm{a}_{1}}+t^{\prime}(e^{i\bm{k}\cdot\bm{a}_{2}}+e^{i\bm{k}\cdot\bm{a}_{3}}).$ (15b)
@@
$\int_{\mathrm{BZ}}\frac{d^{d}k}{(2\pi)^{d}}\sum_{n}f_{n\bm{k}}\partial_{k_{i}}m^{\mathrm{orb}}_{n\bm{k}},$ (16)
where $f_{n\bm{k}}=f(\epsilon_{n\bm{k}}-\mu)$ and $m^{\mathrm{orb}}_{n\bm{k}}$ is the orbital magnetic moment oriented in the direction perpendicular to the surface. Without stretching $\mathrm{MoS_{2}}$, orbital magnetic moments (OMMs) with opposite signs exist around the $\mathrm{K}$ and $\mathrm{K}^{\prime}$ points [59]; however, the OMMD vanishes due to $\mathrm{C_{3v}}$ symmetry. When stretching $\mathrm{MoS_{2}}$, the OMMD in the same direction as the stretching direction becomes finite and generates the GME. This discussion is similar to the Berry curvature dipole in the nonlinear Hall effect [60]. The GME in the DC limit has been experimentally observed as the orbital Edelstein effect in stretched $\mathrm{MoS_{2}}$ [61, 62]. Above the frequency corresponding to the optical gap $\sim 2\mu$, interband effects become finite around $\mathrm{K}$ and $\mathrm{K}^{\prime}$ points.
@@
IV.2 peak shift in optical activity
This peak shift is unique to the orbital optical activity in metals. It becomes clear when the optical activity is represented by the Bloch band basis. The full representation is derived in a recent work [27] (see Eq. (26) in the paper). It shows that there is a Fermi surface effect (This effect originates from a term including the derivative of the Fermi distribution function. This term only becomes finite in metals) affecting the interband transition besides the Fermi sea effect (arising from terms which do not include the derivative of the Fermi distribution function. These terms describe transitions between unoccupied and occupied states) as the antisymmetric part of the optical activity $\sigma^{(A)}_{\mu\nu\lambda}(\Omega)$ includes [27]
$M_{\mu\nu\lambda}(\Omega)\equiv$
$\int_{\mathrm{BZ}}\frac{d^{d}k}{(2\pi)^{d}}\sum_{n\neq m}\frac{-\tilde{\Omega}\partial_{k_{\lambda}}f_{n\bm{k}}\epsilon_{mn\bm{k}}}{\epsilon_{mn\bm{k}}^{2}-\tilde{\Omega}^{2}}\mathrm{Im}\Bigl{(}\mathcal{A}^{\mu}_{nm\bm{k}}\mathcal{A}^{\nu}_{mn\bm{k}}\Bigr{)}.$ (18)
Here, we define $\epsilon_{mn\bm{k}}=\epsilon_{m\bm{k}}-\epsilon_{n\bm{k}}$, $\tilde{\Omega}=\Omega+i\Gamma$, and $\mathcal{A}^{\mu}_{nm\bm{k}}=i\braket{u_{n\bm{k}}}{\partial_{k_{\mu}}u_{m\bm{k}}}$ is the Berry connection. This term is unique to the orbital contribution and does not exist in the spin contribution, and it has been missed in previous works. Ref. [27] has derived it, recently. It demonstrates that the effect of optical transitions in the optical activity cannot be simply described by transitions from occupied to unoccupied states, which is described by the Fermi sea term. We note that in this two-band model with particle-hole symmetry, the Fermi sea term is suppressed. This Fermi surface term is dominant and is locally enhanced at the optical gap $2\mu$ because only the states satisfying $\epsilon_{n\bm{k}}\sim\mu$ contribute and only the state satisfying $\epsilon_{m\bm{k}}\sim-\mu$ due to the particle-hole symmetry, which is consistent with the result in Fig. 2.
In addition, we obtain a sum rule related to the shifting peak structure from $M_{\mu\nu\lambda}(\Omega)$. The sum rule reads
[…]
$=\frac{\pi}{2}\varepsilon_{\mu\nu\theta}\int_{\mathrm{BZ}}\frac{d^{d}k}{(2\pi)^{d}}\sum_{n}f_{n\bm{k}}\partial_{k_{\lambda}}m^{\mathrm{orb},\theta}_{n\bm{k}}.$ (19)
Here, we define the orbital magnetic moment $\bm{m}^{\mathrm{orb}}_{n\bm{k}}=\frac{1}{2}\mathrm{Im}\bra{\bm{\nabla}u_{n\bm{k}}}\times(\epsilon_{n\bm{k}}-H_{\bm{k}})\ket{\bm{\nabla}u_{n\bm{k}}}$. This sum rule means that the area of this shifting peak structure is determined by the OMMD.
IV.3 Optical rotation
The optical activity $\sigma_{xyx}$ in our model can be observed using the optical rotation, the phenomenon in which the polarized axis of linearly polarized light rotates. The optical rotation is usually caused by the optical Hall effect in magnets without time-reversal symmetry, which is called the Kerr effect and the Faraday effect. Because the present model obeys time-reversal symmetry, the optical rotation originating from the optical Hall effect does not occur. However, the optical activity also generates an optical rotation by a different mechanism, as described in the following discussion.
The spectrum of the longitudinal optical conductivity $\sigma_{ii}(\omega)$ can be extracted from the experimentally observed complex reflectance. Similarly, the spectrum of the optical activity can be extracted from the experimentally observed rotational angle and the ellipticity caused by the optical rotation. In this subsection, we discuss the optical rotation as a possibility of observing the optical activity experimentally.
@@ V Model calculation in the superconducting state
Figure 5: Numerical results of the rotational angle $\theta$ and ellipticity $\varepsilon$ for the reflected light (Left) and the transmitted light (Right) on a logarithmic scale. We denote the superconducting state and the normal state as sc and n, respectively. We use the same parameters and units as in Fig. 1. We set $\Delta=0.005$.
The optical rotation is also changed because of the spectral modification by the superconducting state. We show the rotational angle $\theta$ and ellipticity $\varepsilon$ in Fig. 5 (blue lines). We also show the results of the normal state for comparison (black lines). As discussed in Sec. IV.3, in the normal state, Eqs. (20) and (21) are almost purely imaginary in the Drude regime. Thus, the phase difference of the electric fields is almost $\pi/2$, and the ellipticity is dominant. In the superconducting state, for frequencies inside the superconducting gap, the dominant components of $\sigma_{xx/yy}$ and $\sigma_{xyx}$ are changed to the imaginary part and the real part, respectively. Therefore, Eq. (20) is no longer purely imaginary, and the rotational angle is enhanced. At low frequencies, $\mathrm{Im}\sigma_{xx}$ increases by $\Omega^{-1}$ and becomes comparable to the inverse vacuum impedance $(\mu_{0}c)^{-1}$. Then, the real part and imaginary part of Eq. (20) are comparable, and the rotational angle and ellipticity are in the same order of magnitude for the reflected light. For transmitted light, Eq. (21) is also no longer purely imaginary and becomes rather real when $(\mu_{0}c)^{-1}>\mathrm{Im}\sigma_{xx}$, thus the rotational angle is dominant. As the frequency decreases further, $\mathrm{Im}\sigma_{xx}$ increases by $\Omega^{-1}$ and becomes comparable to $(\mu_{0}c)^{-1}$. Then, the rotational angle and the ellipticity also become comparable in magnitude.